May rides her bike the same distance that Leah walks. May rides her bike 10 km/h faster than Leah walks. If it takes May 1 h and Leah 3 h to travel that distance, how fast does each travel
Answer by TimothyLamb(4379)

x = speed of leah
y = speed of may
---
y = x + 10
---
s = d/t
d = st
---
leah:
d = xt
d = 3x
---
may:
d = yt
d = 1y
---
equate the distances traveled:
3x = 1y
3x = y
3x = x + 10
2x = 10
x = 5
y = x + 10
y = 5 + 10
y = 15
---
answer:
x = speed of leah = 5 kph
y = speed of may = 15 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php